$2000 is deposited into an account for n years at an interest rate of 16% per annum compounded quarterly. After two years the interest rate changes to 13% per annum compounded monthly. At the end of the term the value of the investment is $4200. Calculate the value of n.
1
Expert's answer
2020-04-22T19:39:18-0400

compound interest formula,

"final\ value=P*(1+i)^k\\"

where, i=interest rate k=number of periods.


compounded quarterly for first two years. Therefore i and k are consider quarterly,

For first two years,

"i=4\%\\k=8 \\p=\$2000"

F1 is the final value after two years,

"F_1=2000(1+0.04)^8"


compounded monthly for next n-2 years. Therefore i and k are consider monthly

For next n-2 years,

"i=\frac{13}{12}\% \\ k=12(n-2)\\p=F_1\\final\ value=\$4200"

"4200=F_1(1+\frac{13}{12}*0.01)^{k}\\
(1+\frac{13}{12}*0.01)^{k}=\frac{4200}{F_1}\\
k*\ln(1+\frac{13}{12}*0.01)=ln(\frac{4200}{2000(1+0.04)^8})\\
k=39.74"


"k=12(n-2)\\
n=\frac{k}{12}+2=\frac{39.74}{12}+2=5.312"

n=5.312years=5 years and 4 months.